OpenStudy (anonymous):

Aaron deposited $591.07 in a savings account that earns 3.7% simple interest. What is Aaron’s account balance after seven years?

OpenStudy (anonymous):

Fromula for finding simple interst http://math.about.com/od/businessmath/ss/Interest.htm

OpenStudy (anonymous):

@signedup : try once

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

744?

OpenStudy (anonymous):

744.16 ?

OpenStudy (anonymous):

I = PRT I = (591.07)(.037)(7) I = 153.087 rounds to 153.09 591.07 + 153.09 = 744.16 I agree with you :)

OpenStudy (anonymous):

excellent thank you

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

do you want to help me more?

OpenStudy (anonymous):

if I can....I have a little time left before I have to go

OpenStudy (anonymous):

@signedup :all are here to help only

OpenStudy (anonymous):

Mark currently has a balance of $973.70 in an account he has held for 17 years. He opened the account with an initial deposit of $700. What is the simple interest on the account

OpenStudy (anonymous):

is it 8.2 ?

OpenStudy (anonymous):

now we are looking for R in the formula. So I am going rearrange the formula. I = PRT : I/PT = R your interest earned was 973.70 - 700 = 273.70 I/PT = R 273.70/(700)(17) = R 273.70/11900 = R .023 = R 23% = R <-- this is your interest rate check... I = PRT 273.70 = (700)(.023)(17) 273.70 = 273.70 (correct)

OpenStudy (anonymous):

wait...are you looking for interest or interest rate ?

OpenStudy (anonymous):

simple interest

OpenStudy (anonymous):

simple interest is 973.70 - 700 = 273.70

OpenStudy (anonymous):

percentage

OpenStudy (anonymous):

if you want percentage, then you are looking for interest rate, which is 23%

OpenStudy (anonymous):

its not an option

OpenStudy (anonymous):

1.7 2.3 5.9 8.2

OpenStudy (anonymous):

%

OpenStudy (anonymous):

these are the options

OpenStudy (anonymous):

I am sorry...I do not know

OpenStudy (anonymous):

ok thanks for trying

OpenStudy (anonymous):

How did you get 8.2 ?

OpenStudy (anonymous):

can you solve this I=P r T 973.70 = 700 x R x17 r = 973.70/700/17

OpenStudy (anonymous):

its steps to answer the question

OpenStudy (anonymous):

someone told me but wasnt sure

OpenStudy (anonymous):

so i just thought i should check with others

OpenStudy (anonymous):

973.70 = (700)(R)(17) 973.70 = 11900R 973.70/11900 = R .0818 rounds to .082 = R 8.2% = R that works.

OpenStudy (anonymous):

good

OpenStudy (anonymous):

thanks for confirming

OpenStudy (anonymous):

sorry about my mess up....it probably confused you.

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

8.2 is wrong